12. A frog hops so that it leaves the ground with a speed of vo and an angle of 0 above the horizontal. Write an expression for the frogs maximum height during this leap. 13. Write an expression for the time the frog is in the air in the previous problem.
Answer #1

a) Vy^2 = uy^2- 2gH

where H = maximum height of the frog

At maximum Height Vy = 0

uy^2 = 2gH

H = uy^2 / 2g

= (Vo*sin\theta)^2 / 2g

13) Time of flight will be double of time taken to reach maximum height

Vy = Uy - gt

0 = Vo* sin \theta - gt

t = Vo sin \theta / g

total time taken = 2t = 2*Vo sin \theta / g